OpenStudy (anonymous):

Get a medal :3 What value of k solves the equation? k ^-2 = 1/64 A. 1/8 B. 8 <--- I think it's this answer >.< C. 32 D. 128

OpenStudy (anonymous):

b

OpenStudy (anonymous):

1/k^2 = 1/64 k^2 = 64 k = 8

OpenStudy (anonymous):

@Rohith15 @Saeeddiscover Ok thank you, just wanted to make sure. Rohith15 gave a more detailed answer, so the medal goes to you :3